After graphing your standard curve for the Bradford Analysis and plotting the Unknown A BSA sample (unknown concentration), you found that it had an absorbance of 0.48. Using the linear regression below, calculate the [BSA].

y= 0.0056x - 0.0047

r2= 0.9938
Answer #1

Answer:

Given the equation of the standard curve is

y= 0.0056x - 0.0047 with r2=0.9938.

where x=[BSA] and y=absorbance.

Given absorbance of unknown BSA sample=0.48

then [BSA]=?

y= 0.0056x - 0.0047

0.48= 0.0056 [BSA] - 0.0047

0.0056 [BSA]=0.48 + 0.0047=0.4847

[BSA]=0.4847/0.0056=86.55

[BSA]=86.55.

Please let me know if you have any doubt. Thanks